OpenStudy (anonymous):

simplify: Cos^2x/(1+sinx)

OpenStudy (anonymous):

using trig identities

OpenStudy (mertsj):

replace cos^2x with 1-sin^2x

OpenStudy (mertsj):

Then factor the numerator and reduce the fraction.

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

ok

OpenStudy (rulnick):

oops said too much sorry

OpenStudy (anonymous):

?

OpenStudy (mertsj):

@ rulnick Let the asker do the problem

OpenStudy (rulnick):

indeed, my apologies

OpenStudy (rulnick):

cos^2=1-sin^2 so cos^2/(1+sin) = (1-sin^2)/(1+sin) but (1-sin^2) can be factored as (1-sin)(1+sin) which allows cancellation.

OpenStudy (anonymous):

oh ok. i understand now. Thanks so much @rulnick !! :)

OpenStudy (rulnick):

very welcome